OpenStudy (anonymous):

what is the slope of the line that passes through points (-6,1) and (4,-4)? A) -2 B) 2 C) -1/2 D) 1/2

OpenStudy (anonymous):

slope/gradient formula is \[m=\frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

m = -4-1 / 4--6 = -5/ 4+6 = -5/10 = -1/2

OpenStudy (anonymous):

so C

OpenStudy (anonymous):

yea thanks i just did some miscalculations